Circular No. 7454

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

EXO 1745−248 = XB 1745−25

C. B. Markwardt, University of Maryland and Goddard Space Flight Center (GSFC); and J. H. Swank, GSFC, report: “RXTE PCA scans of the galactic bulge on July 2.73 and 6.69 UT have detected a transient source. The position is consistent with a reappearance of the source EXO 1745−248 = XB 1745−25, an x-ray burster in the globular cluster Terzan 5 (cf., e.g., IAUC 3506). The best position as observed by ROSAT was α = 17^{h}48^{m}05^{s}.0, δ = −24^{o}46′47′′ (equinox 2000.0; Johnston et al. 1995, A.Ap. 298, L21). The RXTE observations are inconsistent with the position of Johnston’s nearby source S2 (= 1RXS 174746.0−24451; 6′.9 away) at a high confidence level. The 2–10-keV x-ray fluxes in the PCA observations were 54 and 94 mCrab, indicating that the intensity is rising. Pointed PCA observations are being scheduled; follow-up observations by imaging instruments are encouraged.”

SUPERNOVA 2000cu IN ESO 525-G004 S. Benetti, R. M. Dominguez and A. Zacchei, Telescopio Nazionale

Galileo (TNG); E. Molinari, Osservatorio Astronomico di Brera; and E. Giro, Osservatorio Astronomico di Padova, report: “A fully reduced CCD spectrum (range 331.4–721.5 nm, resolution 1.2 nm), obtained on July 14.03 UT with the TNG reflector (+ D.o.lo.res spectrograph), shows SN 2000cu (cf. IAUC 7453) to be a type-Ia supernova at maximum light. The spectrum is dominated by P-Cyg lines of Ca II H and K, S II and Si II ions, where the absorption component of the Si II 597.2-nm transition is very intense for the deduced phase.”

COMETS C/1999 H5, 2000 M9, 2000 N1 Following the item on IAUC 7453, three more Kreutz sungrazing com-

ets have been detected with the SOHO/LASCO coronagraphs (C2 alone for the first, both C3 and C2 for the others). Detections were by M. Oates (in archival data), by M. Meyer, and by T. Lovejoy and M. Boschat, respectively. Discovery positions follow. Details are on MPEC 2000-N34.

1999 UT  α_{2000}  δ_{2000}  Comet
Apr. 24.455  2^{h}12.^{m}5  +11^{o}43′  C/1999 H5
2000 UT  α_{2000}  δ_{2000}  Comet
June 30.971  6^{h}31.^{m}6  +20^{o}16′  C/2000 M9
July 1.929  6 34.8  +19 24  C/2000 N1
2000 July 14  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Brian G. Marsden